OpenStudy (anonymous):

Can someone help me find the Inverse Laplace Transform of L^-1{(e^-2s)/(s^2+s-2)}

OpenStudy (anonymous):

I'll try, give me a minute to review laplace transforms, been a couple of semesters since diffeq. I can tell you right off the bat though, that you will need a step function, and will need to split the bottom up into partial fractions.

OpenStudy (anonymous):

hmm Okay... the e term on the top is the part that confuses me.

OpenStudy (anonymous):

See it like this: \[e^{-2s}\frac{1}{s^{2}+s-2}\]

OpenStudy (anonymous):

Ok, give me a minute, I am going to attempt to solve it to make sure I understand it correctly. Thomas is correct, that is how you have to look at it. Then you have a function like: \[\frac{1}{(s+2)}*...\]

OpenStudy (anonymous):

\[e^{-2s}\frac{1}{s^{2}+s-2}=e^{-2s}\frac{1}{(s+2)(s-2)}\]

OpenStudy (anonymous):

@Thomas, slight typo, should be (s+2)(s-1)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Wait, made a mistake there.

OpenStudy (anonymous):

\[A/(s-2)+b/(s-1)=e ^{-2s}\] or \[A /(s-2)+b/(s-1)=1\]

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

Then worry about the e afterwards?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The power of e is just a shift in the domain, so that's easiest to do last.

OpenStudy (anonymous):

so \[A(s-2)+B(s-1)=1\] s=2 => B=1 s=1=> A=-1 Which will make it: \[L ^{-1}\left\{ e ^{-2s}\left(( -1/(s-1))+(1/(s-2) \right) \right\}\] and then just solve like this

OpenStudy (anonymous):

it's A(s+2)+B(s-1)=1

OpenStudy (anonymous):

But your approach is good.